# Towards Understanding Multi-Task Learning (Generalization) of LLMs via Detecting and Exploring Task-Specific Neurons

### v1

@@ A.4 Results of Identifying Neurons Experiments in Bloom-7b1
A.5 Training-Test Task Combination
Table 6 shows the training and test task combinations used to investigate generalization and specialization.
A.6 Ablation Experiments
@@ A.7 Overlap Rate
Table 8 shows the overlap rate of task-specific neurons between the training tasks and test tasks. It is worth noting that for all training-test task combinations listed in Table 6, we use the overall set of task-specific neurons of three training tasks as $\mathcal{N}_{x}$ and the overall set of task-specific neurons of three test tasks as $\mathcal{N}_{y}$.
A.8 Results of Correlation Coefficients
Table 7 shows the correlation coefficients between the similarity of task-specific neuron parameters and the generalization performance.
A.9 Benchmarks of Continuous Learning
@@ A.10 Task Orders
A.11 Metrics for Catastrophic Forgetting Experiments
Let $a_{i,j}$ be the testing accuracy of the $i$-th task after training on $j$-th task, and $A_{i}$ denote the testing accuracy after training on task $i$ alone. The evaluation metrics are:
    $\mathrm{CL}=\frac{1}{N}\sum_{i=1}^{N}a_{i,N}$ (12)
A.12 Baselines for Catastrophic Forgetting Experiments
We used the following continual learning techniques as baselines:
  • SeqFT: training the entire model parameters on a sequence of tasks.
  • SeqLoRA: training fixed-size LoRA parameters on a sequence of tasks.
  • Per-Task FT: training a separate model for each task.
@@ A.12 Baselines for Catastrophic Forgetting Experiments
Group Training Tasks ID Test Tasks OOD Test Tasks
(a) Amazon, QQP, MNLI Amazon, QQP, MNLI SST-2, Paws, GPTNLI
Tweetqa, CommonGen, Xsum
(b) Sciqa, E2E, CNN Sciqa, E2E, CNN SST-2, Paws, GPTNLI
Tweetqa, CommonGen, Xsum
@@
Testset SST-2 Paws GPTNLI Tweetqa CommonGen Xsum
r p-value r p-value r p-value r p-value r p-value r p-value
PCCs 0.87 0.02 0.92 0.01 0.79 0.05 0.96 0.00 0.96 0.00 0.97 0.00
SROCC 0.81 0.05 0.77 0.07 0.81 0.05 0.77 0.07 0.83 0.04 0.71 0.11
@@
Table 8: The overlap rate of task-specific neurons between training tasks and test tasks when controlling the proportion of task-specific neurons. The overlap rate formula can be found in Appendix A.7.

### v2

@@ A.4 Additional Experiments for Identifying Neurons
Figure 5 (a) and (b) show the performance on all tasks after deactivating a particular class of task-specific neurons for Llama-2-7b on six classification and six generation tasks, respectively. In both $6\times 6$ matrices, the values on the main diagonal are significantly higher than those at other locations in the same row and column. This suggests that (1) the impact of deactivating a particular class of task-specific neurons on all other tasks is weaker than the impact on this task itself, (2) the impact of deactivating a particular class of task-specific neurons on this task itself is stronger than the impact of deactivating other classes of task-specific neurons on this task.
Figure 5 (c) and (d) show the performance on all tasks after fine-tuning a particular class of task-specific neurons on six classification and six generation tasks for Llama-2-7b, respectively. In both $6\times 6$ matrices, most of the values on the main diagonal are also significantly higher than those of the other elements in the same row and column. The few exceptions are the AmazonFood and SST-2 tasks, which are relatively simple and where zero-shot learning works well enough so that there is little space for improvement. There’s also E2E and CommonGen, which are limited by the difficulty of tasks and have limited scope for improvement. But in each column of these matrices, the impact of deactivating a particular class of task-specific neurons on this task itself is stronger than the impact of deactivating other classes of task-specific neurons on this task.
These results are sufficient to show that our experiments eliminate noise among task-specific neurons and ensure the task-specificity of the neurons we identify.
A.5 Ablation Experiments for Single-task Training
@@ A.6 Overlap Rate
Table 8 shows the overlap rate of task-specific neurons between the training tasks and test tasks. It is worth noting that for all training-test task combinations, we use the overall set of task-specific neurons of three training tasks as $\mathcal{N}_{x}$ and the overall set of task-specific neurons of three test tasks as $\mathcal{N}_{y}$.
A.7 Ablation Experiments on Overlap Rates and Fine-tuning Proportions
Specifically, we chose the fine-tuning neuron proportions as $10\%$, $30\%$, and $50\%$. Under each proportion, we set multiple overlap rates of trained neurons and test task neurons. For example, at a fine-tuning neuron proportion of $10\%$, we first calculate the total number of neurons that need to be trained at this time. We then divide the trained neurons into two sets, a set of task-specific neurons for the test task, and another set containing all remaining neurons. According to the preset overlap rate and the total number of trained neurons, we are able to calculate the number of neurons to be selected from these two sets, and we randomly select them in each of the two sets.
Tables 9, 10 and 11 show the results of the three sets of experiments for classification - classification, generation - generation, and generation - classification, respectively. It can be found that when the proportion of trained neurons is fixed, the performance is improving as the overlap rate increases, which directly proves the conclusion of our paper. In addition to this, when the overlap rate is fixed, the performance is improving as the total number of trained neurons increases. This can be interpreted as a gain from an increase in the number of trained parameters. It is worth noting that when the overlap rate is not fixed, the performance may not be as good as training a small number of neurons despite training more neurons. For example, in the classification - classification experiment, with $30\%$ of trained neurons and an overlap rate of $10\%$, the performance is $81.4$. However, with $10\%$ of trained neurons and an overlap rate of $70\%$, the performance is $82.0$. In all three sets of experiments, the above conclusions hold.
A.8 Benchmarks of Continuous Learning
@@
(a) Deactivation on classification tasks
(b) Deactivation on generation tasks
(c) Fine-tuning on classification tasks
(d) Fine-tuning on generation tasks
Figure 5: Performance of Llama-2-7b on all tasks after deactivation or fine-tuning a particular class task-specific neurons. The element in the $i$-th row and $j$-th column is the performance change for task $j$ due to deactivation or fine-tuning of the task $i$ specific neurons.
@@
Table 8: The overlap rate of task-specific neurons between training tasks and test tasks when controlling the proportion of task-specific neurons.
Overlap rate \ Percentage of trained neurons 10% 30% 50%
10% 80.2 81.4 81.8
20.8% 80.7 - -
30% 81.1 82.0 82.3
50% 81.5 82.3 82.8
53.9% - 82.5 -
70% 82.0 82.7 83.0
84.5% - - 83.1
100% 82.2 83.1 83.6
Table 9: Results at different fine-tuned neuron proportions ($10\%$, $30\%$, $50\%$) controlling the overlap rate under the classification - classification combination. Italics indicate the original experimental results and overlap rates.
Overlap rate \ Percentage of trained neurons 10% 30% 50%
10% 31.6 32.1 32.3
21.6% 32.2 - -
30% 32.5 32.9 33.5
50% 32.7 33.4 33.8
52.5% - 33.8 -
70% 32.9 34.0 34.1
82.0% - - 34.9
100% 33.1 34.4 35.1
Table 10: Results at different fine-tuned neuron proportions ($10\%$, $30\%$, $50\%$) controlling the overlap rate under the generation - generation combination. Italics indicate the original experimental results and overlap rates.
Overlap rate \ Percentage of trained neurons 10% 30% 50%
10% 78.2 78.5 78.6
11.8% 78.4 - -
30% 78.8 78.7 79.3
40.2% - 78.7 -
50% 79.0 79.4 79.7
69.3% - - 80.1
70% 79.3 79.6 80.3
100% 79.5 79.9 80.8
Table 11: Results at different fine-tuned neuron proportions ($10\%$, $30\%$, $50\%$) controlling the overlap rate under the generation - classification combination. Italics indicate the original experimental results and overlap rates.